OpenStudy (chloegauv):

Determine the value of K such that when f(x)=x^4+kx^3-3x-5 is divided by x-2, the remainder is -11.

jhonyy9 (jhonyy9):

so you need divide it and the remainder with k make equal with -11 ok. ?

OpenStudy (chloegauv):

@jhonyy9 so I divided and got 8k+16 as the remainder. I don't understand what you mean by making this equal this -11.

jhonyy9 (jhonyy9):

so you need calculi the value of K so than the reminder being equal -11 yes ?

OpenStudy (chloegauv):

8k+16=-11 8k=-11-16 8k=-27 k=-27/8 k=-3.375 ?

jhonyy9 (jhonyy9):

yes this is right sure so do you understand it now ?

OpenStudy (chloegauv):

Yes, thank you very much!

jhonyy9 (jhonyy9):

but i think inside equation you need give it in fraction form - ok ?

jhonyy9 (jhonyy9):

was my pleasure and anytime bye bye good luck

OpenStudy (chloegauv):

oh i see, ok. You too!

jhonyy9 (jhonyy9):

ty